solve for x in the proportion: 3/5=x/25


Thanks everyone for all the help!

Answer by stanbon(75887)

solve for x in the proportion: 3/5=x/25
------
3/5 = x/25
---
Multiply both sides by 25 to get:
x = 25(3/5)
---
x = 5*3
x = 15
==============
cheers,
Stan H.